OpenStudy (anonymous):

41v - 33 = 42v - 97 V = ? @Shay17

OpenStudy (anonymous):

v=64

OpenStudy (anonymous):

41v-33=42v-97 add 33 41v=42v-64 subtract 42v -1v=-64 divide -1 v=64

OpenStudy (anonymous):

Medal ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ohh ^

OpenStudy (anonymous):

Thanks @keiyuchan